OpenStudy (anonymous):

Probability! You roll a 6-sided die and toss a coin. What is the probability of rolling a number less than 3 and getting heads? I also need help with others such as... If you roll 2 dice what is the probability of rolling a total of 10? <-- How do I find the probability for this? I know that 2 dice would be 6x6 = 36 so the probability would be something out of 36. But how do I find the numerator of the fraction?

OpenStudy (anonymous):

P(1 or 2 and H)=(1/3)*(1/2) P(sum of 10) = P(4, 6) + P(5, 5) + P(6, 4) = 3/36 = 1/12

OpenStudy (anonymous):

Question: Why multiply (1/3) and (1/2)?

OpenStudy (anonymous):

rolling a die and flipping a coin are independent events. => P(A and B) = P(A)*P(B)